A vacuum manufacturer has prepared the following cost data for manufacturing one of its engine components based on the annual production of 50,000 units.

DescriptionCost per MonthDirect Materials  $75,000Direct Labor$100,000Total$175,000


In addition, variable factory overhead is applied at $7.50 per unit. Fixed factory overhead is applied at 150% of direct labor cost per unit. The vacuums sell for $150 each. A third party has offered to make the engines for $60 per unit. 75% of fixed factory overhead, which represents executive salaries, rent, depreciation, and taxes, continue regardless of the decision. Should the company make or buy the engines?


1
Expert's answer
2021-02-28T11:38:36-0500
"Q=50000"

"VC=75000+100000+7,5\times50000=550000"

"FC=1.5 \times \frac{100000}{50000}+0.75\times(30000+375000)=333750"

"TC=FC+VC=550000+333750=883750"

"TR=p\times Q=150\times50000=7500000"

If you buy motors from another company, then


"VC=550000+60\times50000=3550000"

Conclusion: motors must be produced independently